# HiFiSeg: High-Frequency Information Enhanced Polyp Segmentation with Global-Local Vision Transformer

### v1

@@ Abstract
polyp segmentation, colonoscopy, deep learning, pyramid vision transformer
@@ I Introduction
Polyps are abnormal growths in the colon and rectum that protrude above the surrounding intestinal mucosa. Colorectal cancer often develops from colonic polyps, particularly adenomatous ones. Therefore, early detection and removal of these polyps are crucial for preventing the progression to colorectal cancer. Colonoscopy is the gold standard in clinical practice for detecting colorectal lesions [1]. However, manual annotation of polyps is not only labor-intensive but also prone to errors. This highlights an increasing demand for automated and precise image segmentation methods to assist doctors in diagnosis.
Deep learning algorithms represented by CNNs have achieved remarkable success in medical image applications such as cardiac, skin lesion, and polyp segmentation. In particular, fully convolutional networks (FCNs), such as UNet [2], SegNet [3], and DeepLab [4], have become the predominant methods in this field. However, due to the limited receptive fields of CNNs, these CNN-based methods have limitations in capturing long-range relationships and global context, which restricts their representational capabilities in capturing shape and structural information crucial for medical image segmentation.
Transformer [5] reflects complex spatial transformations and long-range feature dependencies through the multi-head self-attention (MHSA) mechanism. It has achieved significant success in the field of natural language processing (NLP). In recent years, the Vision Transformer (ViT) [6] has been proposed to apply the Transformer structure to vision tasks, aiming to overcome the limitations of CNNs in image recognition tasks. Unfortunately, despite the ability of Transformers to capture global dependencies, they struggle with image locality and translational invariance, which affects the accurate segmentation of small targets and boundaries. Therefore, constructing a model that retains both local and global features while preserving their respective characteristics remains a challenging task. Recent works have introduced hybrid architectures of Transformers and CNNs, such as TransUnet [7], HiFormer [8] and LeVit-Unet [9]. These architectures combine the locality of CNNs with the long-range relevance of Transformers to encode both global and local features in medical image segmentation. However, although these methods demonstrate better performance in segmentation tasks, they still have limitations, such as insufficient attention to fine-grained features. The shortcoming hinders the model’s ability to accurately identify small targets and locate boundaries, thereby limiting its effectiveness in learning and generalizing in medical image segmentation. As illustrated in Figure 1, the PraNet model is used as an example to highlight the challenges in polyp segmentation.
@@
Inspired by multiscale and multilevel feature modeling approaches [10, 11, 12, 13], we propose a high-frequency information enhanced polyp segmentation called HiFiSeg. The main components of HiFiSeg are pyramid vision transformer(PVT), global-local interaction module(GLIM), and selective aggregation module(SAM). PVT, a lightweight Transformer with a hierarchical structure, serves as the encoder to capture multiscale features. GLIM employs parallel convolutional kernels and pooling operations of various sizes to aggregate global and local information, enabling the extraction of fine-grained features, which is particularly beneficial for localizing small targets. To further reduce computational complexity, GLIM groups the channels and performs parallel convolutions using depthwise separable convolution. SAM enhances boundary features by using high-level semantic information to selectively guide low-level detail refinement. In summary, our contributions in this paper are as follows:
•We propose a novel framework for colon polyp segmentation, named HiFiSeg. The HiFiSeg adapts the pyramid vision transformer as an encoder to extract more robust features than CNN-based methods.
•We design global-local interaction module(GLIM) and selective aggregation module(SAM) to support our framework. GLIM enhances segmentation performance for small targets by extracting high-level local multi-scale information. SAM addresses boundary ambiguity by selectively fusing low-level boundary details with high-level semantic information.
•We conduct a set of experiments on five standard benchmark datasets for polyp segmentation(Kvasir [14], CVC-ClinicDB [15], CVC-300 [16], CVC-ColonDB [17], ETIS [18]) and compare the results with those from state-of-the-art methods.On the challenging CVC-ColonDB and ETIS datasets, the mDice scores achieved 0.826 and 0.822, respectively, surpassing existing state-of-the-art methods.
@@ II-B Vision Transformer
Transformer [5] is a deep learning model proposed by Vaswani et al. that uses a multi-head self-attention mechanism to capture long-range dependencies in input sequences. The Transformer model was initially applied to natural language processing such as machine translation and text generation. Due to its powerful parallel processing and ability to capture long-range dependencies, it is now widely used in image processing, speech recognition, and other fields. The Vision Transformer (ViT) [6], proposed by Dosovitskiy et al., is the first pure Transformer for image classification tasks. ViT divides an image into fixed-size patches, which are then projected and input into the Transformer model for processing and feature learning. Subsequently, Transformer-based pyramid structures applicable to downstream tasks have been proposed, such as Swin Transformer [25], PVT [10], and Segformer [26].
In medical image segmentation, Transfuse [27] combines Transformers and CNNs in a parallel fashion to efficiently capture global dependencies and low-level spatial details in a much shallower way. TransUNet [7] uses Transformer as encoder for medical image segmentation tasks, combined with U-Net to enhance local details by recovering local spatial information. Polyp-PVT [28] uses PVT as an encoder and designs a similarity aggregation module based on graph convolution. ColonFormer [29]models global semantic relations at multiple scales based on transformer and incorporates a new jump-joining technique for refining the boundaries of polyp objects in global maps. DuAT [30] proposes a dual-aggregate Transformer network in which the global-to-local spatial aggregation module is used to locate large and small targets, and the selective boundary aggregation module is used to preserve boundary details and locate recalibrated objects.

### v2

# This work was supported in part by the Science and Technology Innovation Program of Higher Education Institutions in Shanxi Province 2024L386

@@ Abstract
polyp segmentation, colonoscopy, deep learning, pyramid vision transformer.
[…]
=-21pt
@@ I Introduction
Polyps are abnormal growths in the colon and rectum that protrude from the intestinal mucosa. Colorectal cancer frequently arises from colonic polyps, especially adenomatous ones, making early detection and removal critical for preventing cancer progression. Colonoscopy is widely regarded as the gold standard for detecting colorectal lesions [1]. However, the manual annotation of polyps during colonoscopy is both time-consuming and prone to human error, underscoring the need for automated and accurate image segmentation methods to assist in diagnosis.
Deep learning algorithms, particularly CNNs, have achieved significant success in medical image applications such as cardiac, skin lesion, and polyp segmentation [2]. Fully convolutional networks (FCNs), including models like UNet [3], SegNet [4], and DeepLab [5], have become the dominant approaches in this domain. However, due to the limited receptive fields of CNNs, these methods struggle to capture long-range dependencies and global context, which are essential for accurately representing shape and structural information in medical image segmentation.
The transformer [6] architecture, with its multi-head self-attention (MHSA) mechanism, excels at capturing complex spatial transformations and long-range dependencies. While it has seen tremendous success in natural language processing (NLP), its adaptation to vision tasks through the vision transformer (ViT) [7] was aimed at overcoming the limitations of CNNs in image recognition. However, despite Transformers’ ability to model global dependencies, they struggle with capturing image locality and maintaining translational invariance, which is critical for accurately segmenting small targets and boundaries. To address this challenge, recent works have proposed hybrid architectures [8, 9, 10] that combine the strengths of both Transformers and CNNs, such as TransUnet [11], HiFormer [12], and LeVit-Unet [13]. These models aim to leverage the locality of CNNs with the global context captured by Transformers, enabling them to encode both local and global features for medical image segmentation. While these hybrid models have shown improved performance, they still face limitations, particularly in capturing fine-grained details. This shortcoming affects the accurate identification of small targets and boundary localization, hindering the model’s ability to generalize effectively in medical image segmentation. As illustrated in Figure 1, the PraNet model highlights some of the persistent challenges in polyp segmentation, underscoring the need for better methods to address these issues.
@@
Inspired by multiscale and multilevel feature modeling approaches [14, 15, 16, 17], we propose a high-frequency information-enhanced polyp segmentation framework, termed HiFiSeg. The main components of HiFiSeg include the pyramid vision transformer (PVT), the global-local interaction module (GLIM), and the selective aggregation module (SAM). PVT, a lightweight hierarchical Transformer, serves as the encoder to capture multiscale features efficiently. GLIM employs parallel convolutional kernels and pooling operations of varying sizes to aggregate global and local information, allowing the extraction of fine-grained features. This is particularly advantageous for localizing small targets. To reduce computational complexity, GLIM uses grouped channels with depthwise separable convolution. SAM refines boundary features by leveraging high-level semantic information to guide the selective refinement of low-level details.
In summary, our contributions are as follows:
• We propose HiFiSeg, a novel framework for colon polyp segmentation. HiFiSeg utilizes the Pyramid Vision Transformer as an encoder to capture more robust features than CNN-based methods.
• We design two key modules, GLIM and SAM, to enhance the framework. GLIM improves segmentation performance for small targets by extracting multiscale local features, while SAM addresses boundary ambiguity by selectively fusing low-level boundary details with high-level semantic information.
• We evaluate HiFiSeg on five standard benchmark datasets for polyp segmentation, including Kvasir [18], CVC-ClinicDB [19], CVC-300 [20], CVC-ColonDB [21], and ETIS [22]. On the challenging CVC-ColonDB and ETIS datasets, HiFiSeg achieves mDice scores of 0.826 and 0.822, respectively, surpassing existing state-of-the-art methods.
@@ II-B Vision Transformer
Transformer [6], proposed by Vaswani et al., uses multi-head self-attention to capture long-range dependencies. Initially designed for natural language tasks like translation, Transformers are now widely used in image processing and speech recognition due to their parallel processing and global context modeling. Vision transformer (ViT) [7] was the first pure Transformer for image classification, processing images as fixed-size patches. Subsequent models like Swin Transformer [29], PVT [14], and Segformer [30] introduced pyramid structures for improved vision tasks. Meanwhile, diffusion models [31, 32, 33, 34] have become popular for iteratively refining images through noise reduction. When combined with Transformers, they enhance feature extraction and segmentation, improving performance in tasks like medical image segmentation and object detection.
In medical image segmentation, hybrid architectures combining Transformers and CNNs have shown promise. Transfuse [35] integrates Transformers and CNNs to capture global and local features, while TransUNet [11] uses Transformers as encoders and U-Net to refine local details. Polyp-PVT [36] leverages PVT as an encoder with a graph-based similarity aggregation module. ColonFormer [37] models global semantic relations and refines polyp boundaries, while DuAT [38] introduces dual-aggregate Transformers to balance large and small target detection and enhance boundary precision.